OpenStudy (anonymous):

A machine used to make badges costs $150. It costs $2 to make each badge. Part A: Write a function to show the total cost f(n) of making n badges.

OpenStudy (anonymous):

F(x) = 2*x + 150

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

Part B: What is the total cost of making 20 badges?

OpenStudy (anonymous):

40

OpenStudy (anonymous):

substitute x as 20

OpenStudy (anonymous):

Part C: If the cost of the machine was $20 more, what would be the cost of making 7 badges?

OpenStudy (anonymous):

f(7) = (2*(7) + 150) + 20 = 14 + 170 = 184

OpenStudy (anonymous):

thanks so much!

OpenStudy (anonymous):

no worries ahah